OpenStudy (anonymous):

A quadratic equation is shown below: 3x2 - 16x + 2 = 0 Part A: Describe the solution(s) to the equation by just determining the discriminant. Show your work. (3 points) Part B: Solve 9x2 + 3x - 2 = 0 by using an appropriate method. Show the steps of your work, and explain why you chose the method used. (4 points) Part C: Solve 3x2 - 10x + 2 = 0 by using a method different from the one you used in Part B. Show the steps of your work. (3 points)

OpenStudy (jdoe0001):

part A -> http://www.regentsprep.org/Regents/math/algtrig/ATE3/discriminant.htm part B well, used the "appropriate method" then :)

OpenStudy (anonymous):

BUT i dont know the appropriate method

OpenStudy (jdoe0001):

well... have you covered quadratics yet?

OpenStudy (jdoe0001):

these functions are just two quadratic ones....so....is assume you have covered the appropriate methods already

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

a=3 b=-16 c=2

OpenStudy (anonymous):

\[x=\frac{ 16\pm \sqrt{-16^{2}4(3)(2)} }{ 2(3) }\]

OpenStudy (aum):

For part b, factorization will work. 9x^2 + 3x - 2 = 0 9x^2 + 6x - 3x - 2 = 0 3x(3x + 2) - 1(3x+2) = 0 (3x-1)(3x+2) = 0 x = 1/3, -2/3 For part c) use the above formula. 3x^2 - 10x + 2 = 0 a = 3, b = -10, c = 2 \(\Large x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a } = \frac{ 10\pm \sqrt{(-10)^{2}-4(3)(2)} }{ 2(3) } = ?\)

OpenStudy (anonymous):

i dont know how to do that either

OpenStudy (anonymous):

Solution via A - discrimintant of this quadratic equation, step by step: http://www.hackmath.net/en/calculator/quadratic-equation?a=&b=&c=&eq2=3x2+-+16x+%2B+2+%3D+0